# Position Paper: Rethinking Privacy in RL for Sequential Decision-making in the Age of LLMs

### v1

@@ I Introduction
The rise of reinforcement learning (RL) in critical real-world applications [1, 2, 3, 4] has exposed a fundamental tension in AI privacy: How do we protect sensitive information in systems that learn and make decisions over time? Traditional privacy frameworks, built for protecting individual data points in static datasets [5, 6, 7], are increasingly inadequate for modern RL systems where sensitive information exists not just in individual moments but in temporal patterns, behavioral strategies, and collaborative dynamics [8]. These privacy challenges arise directly from RL’s fundamental characteristic of learning through sequential interaction.
@@
While these privacy challenges are fundamental to basic RL systems, the emergence of advanced paradigms has further amplified these concerns, particularly in relation to collaborative dynamics. Federated reinforcement learning (FedRL), where multiple agents share learning experiences while keeping data locally [12, 13, 14, 15, 16, 17, 18], introduces the challenge of protecting not only individual agent data but also emergent collective behavioral patterns. Similarly, large language models (LLMs), such as ChatGPT [19] and DeepSeek [20], refined through reinforcement learning with human feedback (RLHF) [21, 22, 23] extend these collaborative privacy challenges to human-AI interaction, creating additional vulnerabilities around protecting annotator characteristics and cultural information encoded in feedback patterns [24, 25]. Recent analysis has uncovered numerous instances of personal data in publicly available RLHF datasets that had evaded removal [26], highlighting how even carefully curated training data can expose private user information. These advanced paradigms demonstrate how the temporal and behavioral aspects of privacy intertwine with collaborative dynamics, pushing privacy challenges beyond individual agent privacy to encompass group-level patterns and societal concerns.
@@ IV-F Theoretical Bounds on Privacy-Utility Trade-offs
Our Sequential Privacy framework implies fundamental trade-offs between privacy and utility. For behavioral pattern privacy, we can establish:
Theorem IV.1 (Privacy-Utility Trade-off)
For any policy learning mechanism $\mathcal{M}$ satisfying $(\alpha,\beta)$-behavioral privacy, the expected performance gap between the private policy $\pi_{\text{priv}}$ and the optimal non-private policy $\pi^{*}$ is bounded by:
$\mathbb{E}[V^{\pi^{*}}(s_{0})-V^{\pi_{\text{priv}}}(s_{0})]\geq\Omega\left(\frac{1}{\beta}\right)$
This result demonstrates that stronger privacy guarantees (smaller $\beta$) necessarily incur larger performance penalties, quantifying the inherent tension between privacy protection and policy optimization.
@@ VI The Path Forward: A Research Agenda
The preceding sections highlight critical questions for achieving robust privacy in sequential decision-making. Here, we sketch four complementary directions—covering theoretical foundations, mechanism design, collaborative privacy, and practical deployment—that together form a research agenda for sequential privacy in RL.
@@ VII Conclusion
To address these challenges, this paper has introduced the Sequential Privacy framework built on four fundamental principles: multi-scale protection, behavioral pattern protection, collaborative preservation, and context-aware adaptation. This framework represents a paradigm shift from traditional privacy approaches by explicitly addressing the unique temporal, behavioral, and collaborative aspects of reinforcement learning systems. Realizing Sequential Privacy requires innovative theoretical frameworks, mechanisms tailored to temporal correlations, and rigorous evaluations across domains. In practice, tensions between privacy, utility, and interpretability must be carefully navigated, often demanding new technical and ethical paradigms.
@@ Acknowledgment
TODO: to include funding acknowledgment.

### v2

@@ I Introduction
The rise of reinforcement learning (RL) in critical real-world applications [1, 2, 3, 4, 5] has exposed a fundamental tension in AI privacy: How do we protect sensitive information in systems that learn and make decisions over time? Traditional privacy frameworks, built for protecting individual data points in static datasets [6, 7, 8], are increasingly inadequate for modern RL systems where sensitive information exists not just in individual moments but in temporal patterns, behavioral strategies, and collaborative dynamics [9]. These privacy challenges arise directly from RL’s fundamental characteristic of learning through sequential interaction.
@@
While these privacy challenges are fundamental to basic RL systems, the emergence of advanced paradigms has further amplified these concerns, particularly in relation to collaborative dynamics. Federated reinforcement learning (FedRL), where multiple agents share learning experiences while keeping data locally [13, 14, 15, 16, 17, 18, 19, 20], introduces the challenge of protecting not only individual agent data but also emergent collective behavioral patterns. Similarly, large language models (LLMs), such as ChatGPT [21] and DeepSeek [22], refined through reinforcement learning with human feedback (RLHF) [23, 24, 25] extend these collaborative privacy challenges to human-AI interaction, creating additional vulnerabilities around protecting annotator characteristics and cultural information encoded in feedback patterns [26, 27]. Recent analysis has uncovered numerous instances of personal data in publicly available RLHF datasets that had evaded removal [28], highlighting how even carefully curated training data can expose private user information. These advanced paradigms demonstrate how the temporal and behavioral aspects of privacy intertwine with collaborative dynamics, pushing privacy challenges beyond individual agent privacy to encompass group-level patterns and societal concerns.
@@ IV-F Theoretical Bounds on Privacy-Utility Trade-offs
Our Sequential Privacy framework implies that any mechanism satisfying $(\alpha,\beta)$–behavioral privacy must incur a quantifiable performance cost. Below we state a concrete bound under standard finite‑MDP assumptions.
Lemma IV.1 (Privacy–Utility Trade‑off (Sketch))
In any finite MDP with discount $\gamma$, a mechanism satisfying $(\alpha,\beta)$–behavioral privacy (Def. 4.2) must incur
$\mathbb{E}\bigl{[}V^{\pi^{*}}(s_{0})-V^{\pi_{\rm priv}}(s_{0})\bigr{]}\;=\;\Omega\!\bigl{(}(1-\gamma)/\beta\bigr{)}.$
The above results can be obtained by applying the standard performance-difference lemma [52] to the Rényi bound. It mirrors the familiar inverse‑scaling trade‑offs in differentially‐private supervised learning (e.g. $\Omega(1/\epsilon)$ lower bounds in DP‑SGD [53]), but here it applies at the level of entire policies rather than per‑step gradient updates. It concretely demonstrates that any mechanism strongly limiting policy divergence must pay a nontrivial price in expected return.
@@ VI The Path Forward: A Research Agenda
The preceding sections highlight critical questions for achieving robust privacy in sequential decision-making. Here, we sketch four complementary directions that together form a research agenda for sequential privacy in RL.
@@ VII Conclusion
To address these challenges, we introduce the Sequential Privacy framework built on four fundamental principles: multi-scale protection, behavioral pattern protection, collaborative preservation, and context-aware adaptation. Delivering on this vision demands new theory for temporal and group‑level privacy, domain‑aware mechanisms, and standardized evaluations that balance privacy, utility, and interpretability.
@@ Acknowledgment
This research is supported by National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative, the Centre for Frontier Artificial Intelligence Research, Institute of High Performance Computing, A*STAR, and the College of Computing and Data Science at Nanyang Technological University. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author and do not reflect the views of National Research Foundation, Singapore, and Infocomm Media Development Authority.